Hanson, James, and Smith, a partnership, is in the process of liquidating. The partners have the following capital account balances; $48,000, $48,000, and ($18,000) respectively. The partners share all profits and losses 16%, 48%, and 36%, respectively. Smith has indicated that the ($18,000) deficit will be covered with a forthcoming contribution. The remaining partners have requested an immediate distribution of $40,000 in cash that is available. How should this cash be distributed??

            A)    Hanson $10,000; James $30,000.          

            B)    Hanson $34,000; James $6,000.

            C)    Hanson $22,308; James $17,692.

            D)    Hanson $28,594; James $11,406.

            E)    Hanson $25,000; James $15,000.
Answer #1

Hanson

James

Smith

48,000

48,000

-18,000

16%

48%

36%



-18,000

40,000




16/(16+48%)=25.00%

48/(16+48%)=75.00%


48,000+48,000+40,000=136,000



136,000*25%=34,000

136,000*75%=102,000



48,000+48,000=96,000



6,000



answered by: Jack Who